The average size of the fish in a lake is 11.4 inches, with a standard deviation of 3.2 inches. How do I Find the probability of catching a fish longer than 17 inches.

Answer by stanbon(75887)

The average size of the fish in a lake is 11.4 inches, with a standard deviation of 3.2 inches. How do I Find the probability of catching a fish longer than 17 inches.
----------------------------
Find the z-value of 17.
---
z(17) = (17-11.4)/3.2 = 1.75
--------------------------
P(x > 17 in)= P(z > 1.75) = normalcdf(1.75,100) = 0.04
==========================================================
Cheers,
Stan H.